OpenStudy (anonymous):

A triangle has side lengths of 7, 24, and 26. Is the triangle right, acute, or obtuse?

OpenStudy (solomonzelman):

What do you think?

OpenStudy (anonymous):

its right angle aint it

OpenStudy (solomonzelman):

You can check this by \[7^2+24^2=26^2\]if this is true then it's a right angle if not, then it's Not a right angle.

OpenStudy (anonymous):

thank u

OpenStudy (solomonzelman):

But, \[7^2+24^2=\color{red} {25}~~~~~~~~no t~~26\]

OpenStudy (solomonzelman):

So either obtuse or acute, it looks like something similar to this,  ACUTE OR OBTUSE?

OpenStudy (anonymous):

its pbtuse

OpenStudy (anonymous):

obtuse

OpenStudy (solomonzelman):

I would say acute, see why? You can actually prove the triangle is acute by finding the angles. Do you want to do this, it will be short.

OpenStudy (anonymous):

find angles using a scientific calculator!!

OpenStudy (solomonzelman):

You don't have to.

OpenStudy (anonymous):

yes......but still if u don't get the figure!!

OpenStudy (solomonzelman):

you can draw the attitude inside the triangle and easily find the sides, but I think you can see it's acute without doing that, can't you....

OpenStudy (solomonzelman):

I mean easily find the angles, not sides....

OpenStudy (anonymous):

yes.........but because @duckduckduck got the answer as obtuse.I th8 this will help him.:)

OpenStudy (solomonzelman):

Oh, I am back, OK, lets do it.....

OpenStudy (solomonzelman):